OpenStudy (anonymous):

Verify the Identity cos(x+pi/2)=-sinx

zepdrix (zepdrix):

Hey :) `Cosine Angle Sum Identity` will be useful here!

zepdrix (zepdrix):

\[\large\rm \cos\left(\color{royalblue}{A}+\color{orangered}{B}\right)=\cos \color{royalblue}{A} \cos \color{orangered}{B}-\sin \color{royalblue}{A} \sin \color{orangered}{B}\]

zepdrix (zepdrix):

\[\large\rm \cos\left(\color{royalblue}{x}+\color{orangered}{\frac{\pi}{2}}\right)=?\]Plug the values into the formula, and simplify! :)

OpenStudy (anonymous):

cos(x)cos(pi/2)-sin(x)sin(pi/2)?

zepdrix (zepdrix):

Mmm good! :) Now simplify cos(pi/2) and the sin(pi/2)

OpenStudy (anonymous):

-sin(x)

OpenStudy (anonymous):

@zepdrix ?

zepdrix (zepdrix):

You're trying to get to -sin(x), yes. But you're missing a step.\[\large\rm \cos\left(x+\frac{\pi}{2}\right)=\cos x \color{orangered}{\cos\frac{\pi}{2}}-\sin x \color{orangered}{\sin\frac{\pi}{2}}\]You have to come up with values for these orange parts.